The equation y= 5 sin (3x-4) where y is in millimetres,x is in metres and t is in seconds, represent a a wave motion. Determine
a) Frequency (b) Period (c) speed of the wave
1
Expert's answer
2020-10-21T09:52:21-0400

a)


"f=\frac{4}{2\pi}=0.84\ Hz"

b)

"f=\frac{2\pi}{4}=1.57\ s"

c)


"v=(4)(0.005)=0.02\frac{m}{s}"